The partial pressure of oxygen in the lung alveoli is a bit lower than in ambient air, being about 100 mm of mercury, or 0.13 Atm (it is lower than the partial pressure in air mainly because oxygen is continually taken up by the alveolar capillaries and carbon dioxide is continually released into the alveoli). In cell-free blood plasma (or a saline solution formulated to match key characteristics of blood plasma), which lacks red blood cells and therefore lacks hemoglobin, the concentration of oxygen will equilibrate at 37° C at about 0.3 ml O2/100 ml plasma. For whole blood (with hemoglobin), however, the O2 concentration is around 20 ml O2/100 ml whole blood.

By what factor does the presence of hemoglobin enhance oxygen solubility in blood? Show your work.

  1. Given the above, imagine that you are an emergency room physician treating a patient who lost a quarter of his blood in an accident. A paramedic replaced this lost blood with saline solution to keep his blood pressure up. The saline solution contains no hemoglobin since it contains no red blood cells. The patient is short of breath and oxygen levels in his blood are dangerously low. If for some reason you must choose between administering pure (100%) oxygen or giving a transfusion of whole blood to restore the red blood cell count, which would you expect to be more helpful? Address this decision by answering the questions below. Show your work and be as quantitatively explicit as possible.
  1. Under normal conditions, what percentage of blood oxygen content is accounted for by oxygen dissolved in the plasma and what percentage is bound to hemoglobin? By how much do these percentages change after the patient has a quarter of his blood volume replaced with saline solution?
  1. a) By what percentage could blood oxygen content be increased by delivering pure (100%) oxygen instead of allowing the patient to simply breathe ambient air?

b) One obviously can’t deliver oxygen at a concentration higher than 100%, but how else might the partial pressure of the oxygen being delivered be modified to increase the amount diffusing into the blood?

  1. By what percentage could blood oxygen content be increased by transfusing whole blood to restore the red blood cell count? Relative to delivering pure oxygen, is this a more promising or less promising approach for restoring normal blood oxygen levels? Explain.
Answer #1

Oxygen is carried in the blood in two forms:

(1) dissolved in plasma and RBC water (about 2% of the total)

(2) reversibly bound to hemoglobin (about 98% of the total). ...

Thus, at a PO2 of 100 mm Hg (typical value for arterial blood), 100 ml of plasma contains 0.3 ml O2 (or 0.3 vol%).

Henry's Law:

Without hemoglobin: K (M/atm) = C (mol/L, or M) / P (atm)K = 0.003 M / 100 mm of Hg = 0.0231 M / atmWith hemoglobin: K (M/atm) = C (mol/L, M) / P (atm)K = 0.008M / 100 mm of Hg = 0.0615 M / atm

If blood infusion is given to restore the red blood cell count would be more helpful in improving the patient’s health because of Henry’s Law (K = C/P). If administer pure oxygen is administered, it will  only increase the partial pressure of oxygen, while the concentration of hemoglobin in red blood cells remains the same, which wouldn’t be effective. For example, increased the partial pressure of oxygen to 0.21 atm, while keeping the concentration of hemoglobin in blood the same 8mM, oxygen solubility in blood is 0.03 M/atm. Compare that to increase in concentration of hemoglobin in blood 0.001M, the oxygen solubility doubles to 0.06 M/atm.

i) Only 1.5 percent of oxygen in the blood is dissolved directly into the blood itself. Most oxygen—98.5 percent—is bound to a protein called hemoglobin and carried to the tissues.

ii) Breathing 100% oxygen increases the oxygen content of blood by only about 10%. Fully oxygenated hemoglobin carries about 200 ml of oxygen inone liter of blood  and about 24 ml of oxygen/liter of blood per atmosphere canbe carried dissolved in water

iii) Post-trauma patients have an oxygen consumption which is proportional to oxygen delivery, suggesting that tissue oxygen consumption is limited by diffusion. Transfusion of packed red blood cells (RBC), which increases the oxygen-carrying capacity of blood, would be expected to increase mixed venous PO2, thereby improving tissue oxygenation. However, the low P50 of stored blood may increase the affinity of hemoglobin for oxygen and reduce oxygen consumption. To evaluate the net effect of these mechanisms, we studied hemodynamic and oxygen transport parameters before and after RBC transfusion in eight critically ill patients. Mixed venous O2 content was measured directly by fuel cell O2 analyzer, and standard P50 was calculated. Following transfusion of one unit of packed RBC which increased mean hemoglobin from 9.2 +/- 0.3 gm/dl to 10.1 +/- 0.3 gm/dl (p less than 0.01), there were no changes in oxygen delivery (490 +/- 80 ml/min/m2), oxygen consumption (210 +/- 30 ml/min/m2), or mixed venous PO/ (37 +/- 2 Torr). Cardiac index (4.1 +/- 0.71 L/min) decreased by 0.4 L/min/m2 (p less than 0.05). Standard P50 decreased by 4.2 +/- 2.4 Torr following transfusion of two units of RBC (p less than 0.05). Red blood cell transfusion thus failed to increase oxygen consumption in these patients, despite an increase in oxygen content. Thus, RBC transfusion may not improve tissue oxygenation.